# Synergizing Large Language Models and Task-specific Models for Time Series Anomaly Detection

### v3

@@ A.3 Proof of Lemma 1
Therefore, $\mathcal{L}_{\beta}^{*}$ satisfies Assumption 5.1-Assumption 5.3 in (Bu et al., 2024). As proven before, minimizing $\mathcal{L}_{\beta}$ by stochastic gradient descent can be deemed as adding a random noise to the gradient of each sample during the stochastic gradient descent process of $\mathcal{L}_{\beta}^{*}$. Since Assumption 5.1-Assumption 5.3 in (Bu et al., 2024) hold, we can use Theorem 4 in (Bu et al., 2024) to prove that after adding these random noise, $\mathcal{L}_{\beta}$ can also converge as quickly as $\mathcal{L}_{\beta}^{*}$ by $\mathbf{O}(T^{-\frac{1}{4}})$. Besides, $\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})+\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})$. Since $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$, $\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})$. Since $\mathcal{L}_{\beta}$ can converge as proven above, its gradient $\mathbb{E}(\mathbf{g})$ should converge to 00 as iteration step grows up. Thus, as iteration step grows up, $\mathbb{E}(\mathbf{g}_{i,j})$ approaches 0, i.e., $\mathcal{P}$ approaches the one that makes $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$. Since $\mathcal{L}_{\beta}^{*}$ is convex, as long as $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$, $\mathcal{P}$ is the optimal solution of $\mathcal{L}_{\beta}^{*}$. Thus, by using $\mathcal{L}_{\beta}$, $\mathcal{P}$ can converge to the optimal solution of $\mathcal{L}_{\beta}^{*}$ with similar converge rate $\mathbf{O}(T^{-\frac{1}{4}})$, as iteration step approaches infinite.
@@ A.4 Proof of Theorem 3
A.5 Anomaly Types
We use different binary classification standards to categorize the anomalies in Flight 1 and Flight 2. Among them, the most common binary standard is point anomaly and contextual anomaly, which is illustrated in  (Lai et al., 2024). Besides, we also introduce two other kinds of binary classification standards for aircraft monitoring datasets to explore the complementary performance between LLM and TSADM. According to the degree of reliance on expert knowledge for anomaly detection, we can divide anomalies into background anomalies and data pattern anomalies. The background anomalies highly rely on expertise knowledge, while the data pattern anomaly can be identified only by TSADMs. As shown in Fig. 5(a), the valid range of the dependent variable is decided by the independent variable, which should satisfy the laws of physics between different parameters. It is usually hard for a neural network to clearly identify the valid range of the dependent variable, given the independent variable. Thus, in this case, it relies on expert knowledge and belongs to background anomaly. In contrast, in Fig. 5(b), the redundant channel mechanism usually requires that most of the redundant channels can reach a consensus. Thus, different redundant channels should always be similar to each other in normal status. In this situation, the neural network can identify the similarity between different channels without expert knowledge. Thus, it belongs to data pattern anomaly. According to the similarity between anomaly and surrounding time series, we can divide anomalies into gradual anomalies and mutant anomalies. As shown in Fig. 5(c), the anomalies are quite different from surrounding normal time series and they should belong to mutant anomalies. In Fig. 5(d), the anomalies are similar to the surrounding normal time series and they should belong to the gradual anomaly.
@@ A.5 Anomaly Types
A.6 Datasets
In our experiments, we used four datasets from different scenarios: cloud service monitoring, aircraft monitoring, and mathematical assumption-based anomaly datasets. We introduce these datasets in the following.
  • MGAB.The dataset comprises a synthetic Mackey-Glass time series characterized by non-trivial anomalies. Mackey-Glass time series are recognized for their propensity to exhibit chaotic behavior under specific conditions. The MGAB dataset encompasses 10 Megabytes (MG) of time series data, with each series extending to a length of 105 time points. Within each time series, 10 anomalies are meticulously introduced using a procedure outlined subsequently. Unlike other synthetic benchmarks, it poses a considerable challenge for the human eye to discern the artificially introduced anomalies from the inherent chaotic behavior. A segment of a time series featuring 3 anomalies is illustrated in the graph above. Due to the limit of computation resources when using LLM, we only use 10000 time slots in our experiments. Since there are no point anomalies in the original dataset, but our model utilizes the complementary performance of LLM and TSADM on point anomaly and contextual anomaly, we insert some point anomalies into the original dataset and release it in ”Supplementary Material”.
  • Mustang. The Mustang dataset chronicles the task duration of operations over a span of five years. The authors in (Chen et al., 2024b) have conducted preprocessing on the raw dataset, where they count the histogram of the task duration time distribution at each time slot. Subsequently, they manually identified and labeled the slowdown anomalies. Due to the cost of LLM, we only use part of this dataset.
  • Flight 1 & 2. These datasets are collected from an airline company, which is one of the world’s top 150 enterprises. Different data sets are monitoring data of different flight control system components. The datasets use redundant channels (Osder, 1999) to ensure the robustness of flight systems and refer to sophisticated collaborative relationships between different parameters, which introduces complex rules to judge the availability of the flight control system and is hard to learn only by neural networks. Besides, we show the ratios of normality and different kinds of anomaly in Fig. 6, where the anomaly classification is defined in Appendix. A.5.
Figure 6: The distribution of Flight datasets
A.7 Experiment details
@@
$\hat{\lambda}$ The weight of lagrange coefficient in $\mathcal{L}_{a}$
@@
A.8 Architecture of Our Conditional Network
The conditional network uses the data representation obtained from a TSADM as the condition and leverages the aligned anomaly score of the TSADM and anomaly score of LLM to output the collated anomaly score. As shown in Eq. 22, we firstly concat data representation, aligned anomaly score of TSADM and anomaly score of LLM as many conditional networks do (Rakelly et al., 2018). After that, we use multilayer perceptrons to output collated anomaly score as shown in Eq. 23, where $W_{1}$ and $W_{2}$ are trainable parameters, $\sigma_{1}$ is leaky relu, and $\sigma_{2}$ is sigmoid function.
@@ A.8 Architecture of Our Conditional Network
A.9 Architecture of TSADM
The TSADM consists of 6-layer anomaly attention. Each anomaly attention layer is described in Eq. 24, where $G[i,j]=1-\exp^{-\frac{(i-j)^{2}}{\sigma^{2}}}$, $\sigma$ is a learnable parameter and * denotes an element-wise multiplication. This attention mechanism is built on the observation that anomaly could not establish decentralized connection with the whole time series (Xu et al., 2022) and embeds moderate modification compared with the original one. To promote the anomaly attention capacity of dealing with compound periodic time series, different anomaly attention layers are organized by skimming attention mechanism (Chen et al., 2024b).
@@ A.10 Complementary Performance
In CoLLaTe, we mainly utilize the complementary performance of LLM and TSADM on point anomaly and contextual anomaly. Actually, there are other complementary performances between the TSADM and LLM by using another binary classification standard, and we explore them in this section. We introduce two kinds of binary classification standards: the anomalies that need expert knowledge to detect and the anomalies that can be detected only by data patterns, gradual anomalies, and mutual anomalies. These anomalies are illustrated in detail in Appendix. A.5. We divide Flight 2 into pieces and gather the pieces only obtain only one kind of anomaly. We test GPT4 and CoLLaTe on them and compute their average F1 score, which is shown in Fig. 7(b). In Fig. 7(b), we can see that LLM always has better performance on one kind of anomaly and has poorer performance on another in each binary anomaly classification.
(a) GPT4 performance for different distances from examples
(b) The complementary strength of LLM and TSADM
@@
Figure 7: (a) shows the anomaly detection performance of GPT4 when the distance between the designated time slot for anomaly detection and the time slot of the examples increases. (b) The figure shows that KL-divergence between the histogram of LLM anomaly score and aligned TSADM anomaly score decreases as the iteration step grows and compares the KL-divergence between LLM anomaly score and aligned TSADM anomaly score with the KL-divergence between LLM anomaly score and original TSADM anomaly score; (b) The figure shows the F1 score of LLM (GPT4) and TSADM (CoLLaTe) when using different anomaly binary classification criteria.
A.11 Set-up-pitch Prompt
@@
Numerous studies have shown that incorporating similar historical examples into prompts can significantly enhance anomaly detection performance (Liu et al., 2024b). Based on this insight, our prompt is structured into four components: expertise supplementation, task description, input data, and examples. The expertise supplementation component provides the contextual meanings of the various dimensions of the input data within the application domain, along with relevant domain-specific knowledge in professional document, such as typical value ranges and interrelationships between dimensions. The task description component directs the large language model (LLM) to determine whether a given time slot is anomalous and specifies the required output format. The examples component includes both a positive and a negative example to guide the model’s understanding. The detailed prompt design is provided in the following.
@@
While recent works utilize examples from a fixed dataset, many application scenarios exhibit dynamic patterns over time. For instance, in cloud server monitoring, the cloud environment evolves with service updates, deployments, and revocations (Chen et al., 2024a; Ma et al., 2021). Similarly, in aircraft monitoring, normal patterns shift with changes in flight conditions and aircraft states. As illustrated in Fig. 7(a), our experiments with aircraft monitoring data reveal that the performance of GPT-4 degrades as the temporal distance between the anomaly detection time slot and the example time slot increases. To address this, we leverage a TSADM for anomaly detection to periodically label data flow and update the example collection with manual assistance, ensuring that the examples remain relevant and effective over time.
@@
A.12 Min-Max normalization to TSADM anomaly scores

### v4

@@ A.3 Proof of Lemma 1
Therefore, $\mathcal{L}_{\beta}^{*}$ satisfies Assumption 5.1-Assumption 5.3 in Bu et al., (2024). As proven before, minimizing $\mathcal{L}_{\beta}$ by stochastic gradient descent can be deemed as adding a random noise to the gradient of each sample during the stochastic gradient descent process of $\mathcal{L}_{\beta}^{*}$. Since Assumption 5.1-Assumption 5.3 in Bu et al., (2024) hold, we can use Theorem 4 in Bu et al., (2024) to prove that after adding these random noise, $\mathcal{L}_{\beta}$ can also converge as quickly as $\mathcal{L}_{\beta}^{*}$ by $\mathbf{O}(T^{-\frac{1}{4}})$. Besides, we have the equation as shown in Eq. 17.
$\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})+\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})$ (17)
Since $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$, we can obtain $\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})$. Since $\mathcal{L}_{\beta}$ can converge as proven above, its gradient $\mathbb{E}(\mathbf{g})$ should converge to 00 as iteration step grows up. Thus, as iteration step grows up, $\mathbb{E}(\mathbf{g}_{i,j})$ approaches 0, i.e., $\mathcal{P}$ approaches the one that makes $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$. Since $\mathcal{L}_{\beta}^{*}$ is convex, as long as $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$, $\mathcal{P}$ is the optimal solution of $\mathcal{L}_{\beta}^{*}$. Thus, by using $\mathcal{L}_{\beta}$, $\mathcal{P}$ can converge to the optimal solution of $\mathcal{L}_{\beta}^{*}$ with similar converge rate $\mathbf{O}(T^{-\frac{1}{4}})$, as iteration step approaches infinite.
@@ A.4 Proof of Theorem 2
A.5 Train and Inference
During the training process, we use $\mathcal{L}_{a}+\mathcal{L}_{\beta}$ as the loss function. During the inference process, we use the collated anomaly score output by the conditional network as the anomaly score.
A.6 Anomaly Types
We use different binary classification standards to categorize the anomalies in Flight 1 and Flight 2. Among them, the most common binary standard is point anomaly and contextual anomaly, which is illustrated in  Lai et al., (2024). Besides, we also introduce two other kinds of binary classification standards for aircraft monitoring datasets to explore the complementary performance between LLM and TSADM. According to the degree of reliance on expert knowledge for anomaly detection, we can divide anomalies into background anomalies and data pattern anomalies. The background anomalies highly rely on expertise knowledge, while the data pattern anomaly can be identified only by TSADMs. As shown in Fig. 4(a), the valid range of the dependent variable is decided by the independent variable, which should satisfy the laws of physics between different parameters. It is usually hard for a neural network to clearly identify the valid range of the dependent variable, given the independent variable. Thus, in this case, it relies on expert knowledge and belongs to background anomaly. In contrast, in Fig. 4(b), the redundant channel mechanism usually requires that most of the redundant channels can reach a consensus. Thus, different redundant channels should always be similar to each other in normal status. In this situation, the neural network can identify the similarity between different channels without expert knowledge. Thus, it belongs to data pattern anomaly. According to the similarity between anomaly and surrounding time series, we can divide anomalies into gradual anomalies and mutant anomalies. As shown in Fig. 4(c), the anomalies are quite different from surrounding normal time series and they should belong to mutant anomalies. In Fig. 4(d), the anomalies are similar to the surrounding normal time series and they should belong to the gradual anomaly.
@@ A.6 Anomaly Types
A.7 Datasets
In our experiments, we used four datasets from different scenarios: cloud service monitoring, aircraft monitoring, and mathematical assumption-based anomaly datasets. We introduce these datasets in the following. The used datasets of MGAB and Mustang are included in "Supplementary Material". As for Flight 1 and Flight 2, since they refers to business secret, we could not disclose them.
  • MGAB.The dataset comprises a synthetic Mackey-Glass time series characterized by non-trivial anomalies. Mackey-Glass time series are recognized for their propensity to exhibit chaotic behavior under specific conditions. The MGAB dataset encompasses 10 Megabytes (MG) of time series data, with each series extending to a length of 105 time points. Within each time series, 10 anomalies are meticulously introduced using a procedure outlined subsequently. Unlike other synthetic benchmarks, it poses a considerable challenge for the human eye to discern the artificially introduced anomalies from the inherent chaotic behavior. A segment of a time series featuring 3 anomalies is illustrated in the graph above. Due to the limit of computation resources when using LLM, we only use 10000 time slots in our experiments. Since there are no point anomalies in the original dataset, but our model utilizes the complementary performance of LLM and TSADM on point anomaly and contextual anomaly, we insert some point anomalies into the original dataset and release it in "Supplementary Material".
  • Mustang. The Mustang dataset chronicles the task duration of operations over a span of five years. The authors in Chen et al., 2024b have conducted preprocessing on the raw dataset, where they count the histogram of the task duration time distribution at each time slot. Subsequently, they manually identified and labeled the slowdown anomalies. We include used dataset in "Supplementary Material".
  • Flight 1 & 2. These datasets are collected from an airline company, which is one of the world’s top 150 enterprises. Different data sets are monitoring data of different flight control system components. The datasets use redundant channels Osder, (1999) to ensure the robustness of flight systems and refer to sophisticated collaborative relationships between different parameters, which introduces complex rules to judge the availability of the flight control system and is hard to learn only by neural networks. Besides, we show the ratios of normality and different kinds of anomaly in Fig. 5(a), where the anomaly classification is defined in Appendix. A.6.
A.8 Experiment details
@@
$\hat{\lambda}$ The hyperparameter weight in $\mathcal{L}_{a}$
@@
A.9 Architecture of Our Conditional Network
The conditional network uses the data representation obtained from a TSADM as the condition and leverages the aligned anomaly score of the TSADM and anomaly score of LLM to output the collated anomaly score. As shown in Eq. 22, we firstly concat data representation, aligned anomaly score of TSADM and anomaly score of LLM as many conditional networks do Rakelly et al., (2018). After that, we use multilayer perceptrons to output collated anomaly score as shown in Eq. 23, where $W_{1}$ and $W_{2}$ are trainable parameters, $\sigma_{1}$ and $\sigma_{2}$ are activation functions, such as Leaky ReLU.
@@ A.9 Architecture of Our Conditional Network
A.10 Architecture of TSADM
The TSADM consists of 6-layer anomaly attention. Each anomaly attention layer is described in Eq. 24, where $G[i,j]=1-\exp^{-\frac{(i-j)^{2}}{\sigma^{2}}}$, $\sigma$ is a learnable parameter and * denotes an element-wise multiplication. This attention mechanism is built on the observation that anomaly could not establish decentralized connection with the whole time series Xu et al., (2022) and embeds moderate modification compared with the original one. To promote the anomaly attention capacity of dealing with compound periodic time series, different anomaly attention layers are organized by skimming attention mechanism Chen et al., 2024b .
@@ A.11 Complementary Performance
In CoLLaTe, we mainly utilize the complementary performance of LLM and TSADM on point anomaly and contextual anomaly. Actually, there are other complementary performances between the TSADM and LLM by using another binary classification standard, and we explore them in this section. We introduce two kinds of binary classification standards: the anomalies that need expert knowledge to detect and the anomalies that can be detected only by data patterns, gradual anomalies, and mutual anomalies. These anomalies are illustrated in detail in Appendix. A.6. We divide Flight 2 into pieces and gather the pieces only obtain only one kind of anomaly. We test GPT4 and CoLLaTe on them and compute their average F1 score, which is shown in Fig. 5(c). In Fig. 5(c), we can see that LLM always has better performance on one kind of anomaly and has poorer performance on another in each binary anomaly classification.
(a) The distribution of Flight datasets
(b) GPT4 performance for different distances from examples
(c) The complementary strength of LLM and TSADM
(d) Validity of Assumption 3
@@
Figure 5: (a) shows the distribution of flight datasets. (b) shows the anomaly detection performance of GPT4 when the distance between the designated time slot for anomaly detection and the time slot of the examples increases. (c) The figure shows the F1 score of LLM (GPT4) and TSADM (CoLLaTe) when using different anomaly binary classification criteria. (d) We verify that the mould length of $|\mathcal{L}_{\beta}^{*}(\theta_{1})-\mathcal{L}_{\beta}^{*}(\theta_{2})|_{2}$ is less than $L|\theta_{1}-\theta_{2}|_{2}$ and verifies the validity of assumption 3.
A.12 Set-up-pitch Prompt
@@
Numerous studies have shown that incorporating similar historical examples into prompts can significantly enhance anomaly detection performance Liu et al., 2024b . Based on this insight, our prompt is structured into four components: expertise supplementation, task description, input data, and examples. The expertise supplementation component provides the contextual meanings of the various dimensions of the input data within the application domain, along with relevant domain-specific knowledge in professional document, such as typical value ranges and interrelationships between dimensions. The task description component directs the large language model (LLM) to determine whether a given time slot is anomalous and specifies the required output format. The examples component includes both a positive and a negative example to guide the model’s understanding. The detailed prompt design is provided in the following.
@@
While recent works utilize examples from a fixed dataset, many application scenarios exhibit dynamic patterns over time. For instance, in cloud server monitoring, the cloud environment evolves with service updates, deployments, and revocations Chen et al., 2024a ; Ma et al., (2021). Similarly, in aircraft monitoring, normal patterns shift with changes in flight conditions and aircraft states. As illustrated in Fig. 5(b), our experiments with aircraft monitoring data reveal that the performance of GPT-4 degrades as the temporal distance between the anomaly detection time slot and the example time slot increases. To address this, we leverage a TSADM for anomaly detection to periodically label data flow and update the example collection with manual assistance, ensuring that the examples remain relevant and effective over time.
@@
A.13 Min-Max normalization to TSADM anomaly scores